OpenStudy (anonymous):

Solve each equation by graphing x^2+3x-10=0

OpenStudy (anonymous):

x^2 + 3x -10 = (x+5)(x-2) = 0 so solutions are x = -5 and x = 2 

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

y=x^2+3x-10 also note when x=0, y=-10